OpenStudy (anonymous):

Cynthia has a cup full of dimes and quarters that she would like to put in her savings account. When she takes the cup to the bank, the banker tells Cynthia that the 386 coins in the cup are worth $58.10. How many quarters does Cynthia have in her cup? A. 126 B. 130 C. 256 D. 386

OpenStudy (anonymous):

\[ Q + D = 386 \\ D= 386 - Q \\ 25Q + 10( 386 -Q ) = 5810 \\ 25Q + 3860 - 10Q = 5810 \\ 15Q = 5810 - 3860 \\ Q = \frac{1950}{ 15 }\\ Q = 130 \]